# Update on the German and Australasian Optical Ground Station Networks

### v1

@@ 2.2 German Optical Ground Station Network
The DLR manages the integration of several optical ground stations into a German OGS network (GOGSN, cf. Fig. 3): DLR/GSOC is building the Free-space Optical Ground Antenna TAbernas (FOGATA)32 at the Plataforma Solar de Almeria, a solar research site of the Spanish Centre for Energy, Environmental and Technological Research (CIEMAT). DLR/RSC3 (Responsive Space Cluster Competence Center) together with their prime contractor DiGOS Potsdam have implemented the Laser-Bodenstation (engl: Laser Groundstation) Trauen (LaBoT) 33, 34, the SPACE Research Center of the University of the Bundeswehr Munich is currently setting up the affiliated OGS at their Neubiberg (NBB) campus 35 and the Technical University of Berlin is going to realize a combined FSOC and Satellite Laser Ranging (SLR) station at the DLR site Neustrelitz (NSG).
Table 1 gives an overview on the specifications of those ground stations. Network load is handled via a reactive planning process, taking into account spacecraft visibilities and terminal constraints, as well as actual weather conditions at the OGS locations 37. Station telemetry is shared utilizing the upcoming CCSDS standard for Service Management 38 allowing for automated scheduling. In addition, the European Optical Nucleus Network can be accessed via the respective application programming interface (API). All OGSs follow the design of a robotic observatory consisting of an RC-telescope on a precision mount in a slit-type dome and a containerized control section hosting the electronic system. They will be featured with a common instrument (SOFA) to harmonize their air-to-ground interface gearing them towards compatibility with the upcoming CCSDS standard O3K 39.
FOGATA | LaBoT | OGS NBB | OGS NSG (prelim.) | |
Location | Tabernas, Almeria, Spain | Trauen, Munster, Germany | Neubiberg, Munich, Germany | Neustrelitz, Germany |
Min. Satellite Elevation | 10-deg | 20-deg | 5-deg | 5-deg |
Height a.s.l. | 489-m | 72-m | 549-m | 66-m |
Primary Aperture | 600-mm | 700-mm | 700-mm | 700-mm |
Wavelengths (Rx) | 1550-nm, 1064-nm | 1550-nm, 1064-nm | 1550-nm, 1064-nm | 1550-nm, 1064-nm |
Wavelengths (Tx) | 1530-nm, 1590-nm | 1590-nm | 1590-nm | 1590-nm |
Standard Compatibility | CCSDS O3K, SDA | CCSDS O3K | CCSDS O3K | CCSDS O3K |
Mean Availability | 65% | 25% | 34% | 32% |
@@ 2.3 Small Optical Ground Stations Focal-Optics Assembly: SOFA
The DLR’s Small Optical Ground Stations Focal-Optics Assembly (SOFA) 40, shown in Fig. 6, is a small, fully-integrated, low-cost instrument that is to be mounted on COTS or even existing astronomical telescopes enabling optical communication capabilities. As such, a cost-effective and easy deployment of a fully functional optical ground station can be achieved. With its two low-resolution cameras (one each for infrared and day light) the SOFA unit is able to provide all the required information to the telescope mount to achieve a highly accurate open-loop pointing as well as closed-loop tracking of satellites. Detection of the optical signal is realised both with a calibrated InGaAs-PIN photoreceiver for assessment of the optical signal quality and with a free space optical receiver frontend (RFE, small area, fast APD) for data reception. The SOFA unit is equipped with two fiber ports for coupling to an outgoing spatially diverse laser signal to provide a static or modulated beacon to the satellite and for uplink communications.
@@ 3.1 Australian National University
The ANU Quantum Optical Ground Station (QOGS) is a 70 telescope designed for high-performance optical communication. With a small (<25%) central obscuration and no corrector optics the telescope is optimised for high-throughput of visible and infra-red optical signals. QOGS is situated at the Mount Stromlo Observatory in Canberra, Australia in a purpose-built facility including large optics lab fed via Coudé path, in addition to Nasmyth foci for optical instrument hosting. The Coudé room can contain four large optical benches each addressable by optics individually, providing a stable and consistent environment to host both experimental optical instruments and established infrastructure. Multiple locations for optical benches enables equipment from partners, collaborators, and experimental infrastructure under development to access the telescope in an optimum fashion, with only a steering mirror needed to switch between instruments. The telescope is a high-performance PlaneWave Instruments RC700 41 capable of tracking satellites in LEO with a minimum keyhole, even when fully loaded with additional payloads (cite Lafon SPIE2023 124130V). ANU has demonstrated adaptive optics enhanced quantum communication over a horizontal free-space link, and developed adaptive optics for FSOC including satellite-to-ground links 42.
@@ 3.2 University of Western Australia
TeraNet builds on the success of UWA’s demonstrations of free-space optical links for metrology and communications 43, 44, 45, 46, 47. Most notably, through a SmartSat CRC funded project a robust single-mode fibre coupling between a deployable optical terminal on the ground and a drone-mounted retroreflector to establish and maintain a 100 dual-polarisation quadrature phase shift keying (DP-QPSK) link across 700 (1.4 folded length) at LEO tracking rates 48 was demonstrated, followed up with links to both a fixed wing aircraft and a helicopter at 5 line-of-sight distance (10 folded, McCann et al. in prep). Through the Australian Space Agency’s Moon to Mars Feasibility Mission, a NASA-O2O compatible high photon efficiency link to a drone was demonstrated 49.
TeraNet-1, previously dubbed the Western Australian Optical Ground Station 50, is a 0.7 PlaneWave Instruments CDK700 installed on the roof of the physics building at the UWA campus in Perth. It serves as the main research platform where the existing infrastructure and convenient location allow for rapid testing and innovation. First downlink is expected in Q1 2024. TeraNet-2 will be the centerpiece of TeraNet; it will be based on a PlaneWave Instruments RC700 and located at the Mingenew Space Precinct approximately 340 north of Perth. This site experiences one of the lowest fractions of cloud cover in the world 17 and is home to NASA’s MOBLAS-5 Laser Ranging Station, the most productive station in the International Laser Ranging Service network.
@@ 3.4 New Zealand Optical Ground Station Sites
The University of Auckland maintains a field station at Ardmore, in the north of the North Island. This has been the site of a number of radio physics and atmospheric experiments in the past, and is currently home to a number of environmental science projects. The site is flat and at low altitude. It has power, communications, utility services and a basic mechanical workshop facility. Urenui is a settlement on the east coast of the North Island of New Zealand. It is home to the Māori iwi (tribe) Ngāti Mutunga. It lies in a region of lower than average cloud fraction and the iwi facilities were upgraded in 2020 as part of the then Government’s Provincial Growth Fund 51.
In the South Island, the Black Birch Observatory site was used as a centre of operations in an earlier site testing programme for an astronomical observatory52. It was also the southern sky observatory of the US Naval Observatory, operating from around 1982 to around 1995. The facilities on the site were dismantled and removed in 1997 by amateur astronomers and the site was returned to nature. The Mount John Observatory is the home of four professional telescopes: the 1.8 Nishimura telescope, the 1 McLellan telescope, the 0.61 Optical Craftsman and the 0.61 Boller and Chivens telescope, see Fig. 9. Road access is excellent, mechanical and electrical engineering facilities are available onsite, and there is accommodation for extended stays by observers and staff. Birdlings Flat is an area 40 south of Christchurch and has been the location of space-related activities, including launch. Nearby is the the Kaitorete spit, which is the location of the Tāwhaki Joint Venture, a commercial partnership between two rūnanga (a rūnanga is an authority set up by and for a tribe) and the New Zealand Government.
@@
In order to accelerate activity and capability in FSOC operations, New Zealand researchers have developed a prototype FSOC ground station at the Ardmore site. A 10 observatory dome (PD10, Technical Innovations) has been installed and houses an L-350 mount from PlaneWave Instruments, Inc. The intended payload comprises the DLR SOFA (Small Optical Ground Stations Focal-optics Assembly) unit 40 and an RC telescope with a 25 primary. Installation of the OTA and SOFA unit is planned for Q1 2024. For calibration and early testing purposes, a secondary telescope (12 LX200GPS, Meade Instruments) has been mounted on the L-350. A set of three all-sky cameras, developed by NIWA, have been deployed at various sites, including one at Ardmore. These cameras generate cloud image data which is being used in current research into short-term (i.e., on the time-scale of a typical LEO satellite pass) cloud coverage forecasting. Finally, in order to progressively characterise the Ardmore site as an OGS, an integrated seeing monitor (Miratlas) was installed, which comprises a turbulence monitor and weather station.
@@ 4.1 German FSOC Research
Current research areas associated with the FSOC ground stations network in Germany include measurements of the optical channel to analyze and forecast the availability of optical links in dependence on the geographical location. The results will facilitate updates on existing or the definition of new site diversity and channel models. Basic research on the optical channel is the foundation for further developments of FSOC systems. For example, a precise characterization of the optical channel facilitates the development of new transmission techniques such as multiple-input multiple-output (MIMO) for FSO systems 53. A related field of research is the development of new forward error correction codes to counter atmospheric disturbances.
@@ 4.2 Adaptive Optics
FSOC systems use adaptive optics to improve performance 54, 42. The Mount John Observatory (MJO) is one of the leading candidates for an OGS in New Zealand, and as there are existing installed telescopes, work has commenced on developing AO capability at this facility. A low-order (tilt) closed-loop adaptive optics system has been installed on the Boller & Chivens telescope at MJO 55. System design parameters were sourced from earlier work developing a turbulence site profile 56. This system was originally developed to partially correct for the adverse effects of imaging space debris through the Earth’s turbulent atmosphere 57. Target objects included large, defunct satellites, both in LEO and GEO. More recently, the tilt AO corrector was used to improve the stability of fringe patterns generated by a modified Michelson interferometer used to measure spatial coherence of satellites for space domain awareness.
@@ 4.3 Quantum Secured Communications
Quantum key distribution (QKD) is a method in which communicating parties generate and share a secret key where the presence or information leakage to an eavesdropper can be inferred by fundamental quantum mechanics 58. The secret key can then be used to encrypt information sent between parties. QKD networks have been developed globally where the distribution of secret keys have predominantly been fibre-based using DVQKD protocols 59. The most significant demonstration of QKD has been the demonstration of the DVQKD protocol, BB84, from the Micius satellite which distributed secret keys between two optical ground stations in Xinglong, China and Graz, Vienna, 7,600 apart 60. The demonstration achieved a sifted key rate between 3–9 in good weather conditions between a free-space link distance of approximately 1000 and 600 , respectively, validating the feasibility of QKD between LEO satellites and optical ground stations.
In Europe, the European Quantum Communication Infrastructure (EuroQCI) Initiative will make use of innovative quantum communication technologies that are developed, for example, in EU-funded projects such as the Horizon 2020 OpenQKD project 61 or the Quantum Technologies Flagship 62. It will be, furthermore, an integral part of IRIS, the new EU space-based secure satellite communication system 63.
Each optical ground station discussed in this work will in principle be compatible with QKD though a ground stations suitability for a particular mission would need to be assessed based on the compatibility of transmit and receive apertures. For example a 30 transmission aperture on a LEO satellite would require a 1 receive aperture to achieve a positive secret key rate using continuous variable (CV) QKD 64, 65.
@@ 4.4 Network Expansion
A natural expansion of the proposed network would be to include the Antarctic continent in order to provide the continent with a high-bandwidth solution that it currently lacks. Australia and New-Zealand are active in research in Antarctica, occupying coastal locations including Casey, Mawson Davis, and Scott bases. These provide possible locations to start expanding the network to the Antarctic region thanks to the existing logistics and relative ease of access. The geographic location of the continent is advantageous to this technology 6 as polar-crossing orbits are favoured by many satellite operators, and therefore there could be many overhead passes of a polar station per day. The interior of the continent, where internet access is the most challenging, has ideal weather conditions for such technology. The skies are amongst the clearest in the world, with 20 cloud cover 66. There is also ample evidence that seeing conditions are excellent which attracts astronomers to the continent 67, 68, 69. Even during the daytime, which can last for months, there are periods where the seeing drops to less than 1” due to the thermal balance between the atmosphere and the highly radiating ice.
@@ Acknowledgments
NJR contributed to Sections 1, 3.4 and 5. JA contributed to Section 3.4. TT contributed to Section 4.4. MB contributed to Sections 1 and 3. JEC contributed to Sections 3.4 and 4.4. M Sayat and OT contributed to Section 4.3. RTS and AK contributed to Sections 2.2 and 4.1, KM contributed to Section 3.3. DG, JR, MTK contributed to Sections 2 and 4.1. SR contributed to Section 2.2.

### v2

@@ 2.2 German Optical Ground Station Network
The DLR manages the integration of several optical ground stations into a German OGS network (GOGSN, cf. Fig. 3): DLR/GSOC is building the Free-space Optical Ground Antenna TAbernas (FOGATA)32 at the Plataforma Solar de Almeria, a solar research site of the Spanish Centre for Energy, Environmental and Technological Research (CIEMAT). DLR/RSC3 (Responsive Space Cluster Competence Center) together with their prime contractor DiGOS Potsdam are implementing the Laser-Bodenstation (engl: Laser Groundstation) Trauen (LaBoT) 33, 34, the SPACE Research Center of the University of the Bundeswehr Munich is currently setting up the affiliated OGS at their Neubiberg (NBB) campus 35 and the Technical University of Berlin is going to realize a combined FSOC and Satellite Laser Ranging (SLR) station at the DLR site Neustrelitz (NSG). The new Optical Ground Station Oberpfaffenhofen (OGSOP-NG) of DLR/KN is already operational and is used for various FSOC experiments 36.
Table 1 gives an overview on the specifications of those ground stations. Network load is handled via a reactive planning process, taking into account spacecraft visibilities and terminal constraints, as well as actual weather conditions at the OGS locations 38. Station telemetry is shared utilizing the upcoming CCSDS standard for Service Management 39 allowing for automated scheduling. In addition, the European Optical Nucleus Network can be accessed via the respective application programming interface (API). All OGSs follow the design of a robotic observatory consisting of an RC-telescope on a precision mount within a dome and with a (containerized) control section hosting the electronic system. They will be featured with a common instrument (SOFA) to harmonize their air-to-ground interface gearing them towards compatibility with the upcoming CCSDS standard O3K 40.
FOGATA | LaBoT | OGS NBB | OGS NSG (prelim.) | OGSOP-NG | |
Location | Almeria, Spain | Trauen, Germany | Neubiberg, Germany | Neustrelitz, Germany | Oberpfaffenhofen, Germany |
Min. Satellite Elevation | 10-deg | 20-deg | 5-deg | 5-deg | 5-deg |
Height a.s.l. | 489-m | 72-m | 549-m | 66-m | 615m |
Primary Aperture | 600-mm | 700-mm | 700-mm | 700-mm | 800-mm |
Wavelengths (Rx) [nm] | 1064, 1550 | 1064, 1550 | 1064, 1550 | 1064, 1550 | 589, 850, 1064, 1550 |
Wavelengths (Tx) [nm] | 1530, 1590 | 1590 | 1590 | 1590 | 1064, 1590 |
Standard Compatibility | CCSDS O3K, SDA | CCSDS O3K | CCSDS O3K | CCSDS O3K | CCSDS O3K |
Mean Availability | 65% | 25% | 34% | 32% | 34% |
@@ 2.3 Small Optical Ground Stations Focal-Optics Assembly: SOFA
The DLR’s Small Optical Ground Stations Focal-Optics Assembly (SOFA) 41, shown in Fig. 6, is a small, fully-integrated, low-cost instrument that is to be mounted on COTS or even existing astronomical telescopes enabling optical communication capabilities. As such, a cost-effective and easy deployment of a fully functional optical ground station can be achieved. With its two low-resolution cameras (one each for infrared and day light) the SOFA unit is able to provide all the required information to the telescope mount to achieve a highly accurate open-loop pointing as well as closed-loop tracking of satellites. Detection of the optical signal is realised both with a calibrated InGaAs-PIN photoreceiver for assessment of the optical signal quality and with a free space optical receiver frontend (RFE, small area, fast APD) for data reception. The SOFA unit is equipped with two fiber ports for coupling to an outgoing spatially diverse laser signal to provide a static or modulated beacon to the satellite and for uplink communications.
@@ 3.1 Australian National University
The ANU Quantum Optical Ground Station (QOGS) is a 70 telescope designed for high-performance optical communication. With a small (<25%) central obscuration and no corrector optics the telescope is optimised for high-throughput of visible and infra-red optical signals. QOGS is situated at the Mount Stromlo Observatory in Canberra, Australia in a purpose-built facility including large optics lab fed via Coudé path, in addition to Nasmyth foci for optical instrument hosting. The Coudé room can contain four large optical benches each addressable by optics individually, providing a stable and consistent environment to host both experimental optical instruments and established infrastructure. Multiple locations for optical benches enables equipment from partners, collaborators, and experimental infrastructure under development to access the telescope in an optimum fashion, with only a steering mirror needed to switch between instruments. The telescope is a high-performance PlaneWave Instruments RC700 42 capable of tracking satellites in LEO with a minimum keyhole, even when fully loaded with additional payloads (cite Lafon SPIE2023 124130V). ANU has demonstrated adaptive optics enhanced quantum communication over a horizontal free-space link, and developed adaptive optics for FSOC including satellite-to-ground links 43.
@@ 3.2 University of Western Australia
TeraNet builds on the success of UWA’s demonstrations of free-space optical links for metrology and communications 44, 45, 46, 47, 48. Most notably, through a SmartSat CRC funded project a robust single-mode fibre coupling between a deployable optical terminal on the ground and a drone-mounted retroreflector to establish and maintain a 100 dual-polarisation quadrature phase shift keying (DP-QPSK) link across 700 (1.4 folded length) at LEO tracking rates 49 was demonstrated, followed up with links to both a fixed wing aircraft and a helicopter at 5 line-of-sight distance (10 folded, McCann et al. in prep). Through the Australian Space Agency’s Moon to Mars Feasibility Mission, a NASA-O2O compatible high photon efficiency link to a drone was demonstrated 50.
TeraNet-1, previously dubbed the Western Australian Optical Ground Station 51, is a 0.7 PlaneWave Instruments CDK700 installed on the roof of the physics building at the UWA campus in Perth. It serves as the main research platform where the existing infrastructure and convenient location allow for rapid testing and innovation. First downlink is expected in Q1 2024. TeraNet-2 will be the centerpiece of TeraNet; it will be based on a PlaneWave Instruments RC700 and located at the Mingenew Space Precinct approximately 340 north of Perth. This site experiences one of the lowest fractions of cloud cover in the world 17 and is home to NASA’s MOBLAS-5 Laser Ranging Station, the most productive station in the International Laser Ranging Service network.
@@ 3.4 New Zealand Optical Ground Station Sites
The University of Auckland maintains a field station at Ardmore, in the north of the North Island. This has been the site of a number of radio physics and atmospheric experiments in the past, and is currently home to a number of environmental science projects. The site is flat and at low altitude. It has power, communications, utility services and a basic mechanical workshop facility. Urenui is a settlement on the east coast of the North Island of New Zealand. It is home to the Māori iwi (tribe) Ngāti Mutunga. It lies in a region of lower than average cloud fraction and the iwi facilities were upgraded in 2020 as part of the then Government’s Provincial Growth Fund 52.
In the South Island, the Black Birch Observatory site was used as a centre of operations in an earlier site testing programme for an astronomical observatory53. It was also the southern sky observatory of the US Naval Observatory, operating from around 1982 to around 1995. The facilities on the site were dismantled and removed in 1997 by amateur astronomers and the site was returned to nature. The Mount John Observatory is the home of four professional telescopes: the 1.8 Nishimura telescope, the 1 McLellan telescope, the 0.61 Optical Craftsman and the 0.61 Boller and Chivens telescope, see Fig. 9. Road access is excellent, mechanical and electrical engineering facilities are available onsite, and there is accommodation for extended stays by observers and staff. Birdlings Flat is an area 40 south of Christchurch and has been the location of space-related activities, including launch. Nearby is the the Kaitorete spit, which is the location of the Tāwhaki Joint Venture, a commercial partnership between two rūnanga (a rūnanga is an authority set up by and for a tribe) and the New Zealand Government.
@@
In order to accelerate activity and capability in FSOC operations, New Zealand researchers have developed a prototype FSOC ground station at the Ardmore site. A 10 observatory dome (PD10, Technical Innovations) has been installed and houses an L-350 mount from PlaneWave Instruments, Inc. The intended payload comprises the DLR SOFA (Small Optical Ground Stations Focal-optics Assembly) unit 41 and an RC telescope with a 25 primary. Installation of the OTA and SOFA unit is planned for Q1 2024. For calibration and early testing purposes, a secondary telescope (12 LX200GPS, Meade Instruments) has been mounted on the L-350. A set of three all-sky cameras, developed by NIWA, have been deployed at various sites, including one at Ardmore. These cameras generate cloud image data which is being used in current research into short-term (i.e., on the time-scale of a typical LEO satellite pass) cloud coverage forecasting. Finally, in order to progressively characterise the Ardmore site as an OGS, an integrated seeing monitor (Miratlas) was installed, which comprises a turbulence monitor and weather station.
@@ 4.1 German FSOC Research
Current research areas associated with the FSOC ground stations network in Germany include measurements of the optical channel to analyze and forecast the availability of optical links in dependence on the geographical location. The results will facilitate updates on existing or the definition of new site diversity and channel models. Basic research on the optical channel is the foundation for further developments of FSOC systems. For example, a precise characterization of the optical channel facilitates the development of new transmission techniques such as multiple-input multiple-output (MIMO) for FSO systems 54. A related field of research is the development of new forward error correction codes to counter atmospheric disturbances.
@@ 4.2 Adaptive Optics
FSOC systems use adaptive optics to improve performance 55, 43. The Mount John Observatory (MJO) is one of the leading candidates for an OGS in New Zealand, and as there are existing installed telescopes, work has commenced on developing AO capability at this facility. A low-order (tilt) closed-loop adaptive optics system has been installed on the Boller & Chivens telescope at MJO 56. System design parameters were sourced from earlier work developing a turbulence site profile 57. This system was originally developed to partially correct for the adverse effects of imaging space debris through the Earth’s turbulent atmosphere 58. Target objects included large, defunct satellites, both in LEO and GEO. More recently, the tilt AO corrector was used to improve the stability of fringe patterns generated by a modified Michelson interferometer used to measure spatial coherence of satellites for space domain awareness.
@@ 4.3 Quantum Secured Communications
Quantum key distribution (QKD) is a method in which communicating parties generate and share a secret key where the presence or information leakage to an eavesdropper can be inferred by fundamental quantum mechanics 59. The secret key can then be used to encrypt information sent between parties. QKD networks have been developed globally where the distribution of secret keys have predominantly been fibre-based using DVQKD protocols 60. The most significant demonstration of QKD has been the demonstration of the DVQKD protocol, BB84, from the Micius satellite which distributed secret keys between two optical ground stations in Xinglong, China and Graz, Vienna, 7,600 apart 61. The demonstration achieved a sifted key rate between 3–9 in good weather conditions between a free-space link distance of approximately 1000 and 600 , respectively, validating the feasibility of QKD between LEO satellites and optical ground stations.
In Europe, the European Quantum Communication Infrastructure (EuroQCI) Initiative will make use of innovative quantum communication technologies that are developed, for example, in EU-funded projects such as the Horizon 2020 OpenQKD project 62 or the Quantum Technologies Flagship 63. It will be, furthermore, an integral part of IRIS, the new EU space-based secure satellite communication system 64.
Each optical ground station discussed in this work will in principle be compatible with QKD though a ground stations suitability for a particular mission would need to be assessed based on the compatibility of transmit and receive apertures. For example a 30 transmission aperture on a LEO satellite would require a 1 receive aperture to achieve a positive secret key rate using continuous variable (CV) QKD 65, 66.
@@ 4.4 Network Expansion
A natural expansion of the proposed network would be to include the Antarctic continent in order to provide the continent with a high-bandwidth solution that it currently lacks. Australia and New-Zealand are active in research in Antarctica, occupying coastal locations including Casey, Mawson Davis, and Scott bases. These provide possible locations to start expanding the network to the Antarctic region thanks to the existing logistics and relative ease of access. The geographic location of the continent is advantageous to this technology 6 as polar-crossing orbits are favoured by many satellite operators, and therefore there could be many overhead passes of a polar station per day. The interior of the continent, where internet access is the most challenging, has ideal weather conditions for such technology. The skies are amongst the clearest in the world, with 20 cloud cover 67. There is also ample evidence that seeing conditions are excellent which attracts astronomers to the continent 68, 69, 70. Even during the daytime, which can last for months, there are periods where the seeing drops to less than 1” due to the thermal balance between the atmosphere and the highly radiating ice.
@@ Acknowledgments
NJR contributed to Sections 1, 3.4 and 5. JA contributed to Section 3.4. TT contributed to Section 4.4. MB contributed to Sections 1 and 3. JEC contributed to Sections 3.4 and 4.4. MTS and OT contributed to Section 4.3. RTS and AK contributed to Sections 2.2 and 4.1, KM contributed to Section 3.3. DG, JR, MTK contributed to Sections 2 and 4.1. SR contributed to Section 2.2.
@@
The TeraNet project is funded by the Australian Space Agency’s Moon to Mars Demonstrator Mission, the Government of Western Australia, and the University of Western Australia.